OpenStudy (anonymous):

Solve using the addition principle t-7=18

OpenStudy (anonymous):

okay. lets see, t -7 =18 what should you add to t-7 to get t+0?

OpenStudy (anonymous):

in other words, what do you add to -7 to make it 0?

OpenStudy (anonymous):

u set it up like t-7 +7

OpenStudy (anonymous):

okay very good. so if you add 7 to the left hand side of the equation, you must add it to the right hand side also to make the equation equal. so t - 7+7 = 18+7

OpenStudy (anonymous):

so the answer will be T=25

OpenStudy (anonymous):

yes. that is right.

OpenStudy (anonymous):

now what if a problem is 4=b-4

OpenStudy (anonymous):

same thing

OpenStudy (anonymous):

use the same principle. exactly. it is the same thing

OpenStudy (anonymous):

in the earlier problem, you found out t = 25.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so this set up will be like 4+4 = b-4+4

OpenStudy (anonymous):

your equation was t - 7 = 18 so what is t?. if you put t = 25, you get 25 minus 7 is 18. which is exactly right. similarly in your new problem, it is: some number b - 4 gives 4

OpenStudy (anonymous):

right

OpenStudy (anonymous):

the answer will be 8 = b

OpenStudy (anonymous):

just put it in words. b-4 = 4 means " the number BEE minus the number 4, gives me a number 4". so what is BEE?

OpenStudy (anonymous):

exactly. and you know b = 8 because 8 minus 4 is equal to 4.

OpenStudy (anonymous):

oh k thank you

OpenStudy (anonymous):

you are welcome